Startup Night 3 - 20140901 UT

Observers: JHill (LBTO), OKuhn, BRothberg (Tucson)
Telescope Operator: SAllanson (LBTO)
AO Support: JCGuerra
Telescope Support: KNewton
Software Support: DSummers
Instrument Support: JPower (LBTO)
Engineering Support: DAshby

Summary

It was a great night. We started with binocular LBC sky flats, and then some wake-up of LBC-Red along with Blue.

We switched to MODS1 plus LBC-Red in pseudomono mode for the rest of the night. Barry and Olga worked on MODS1 wake-up, calibration, testing and science. Seeing on the guider varied from 0.7 to 1.9 arcsec through the night.

After fighting some LBC-Red power problems (Jenny found the loose connection the following afternoon.), we spent the last 4 hours testing WRS on LBC-Red. Marco Stangalini got to run it live from Rome. Both John and Marco gained significant insight into how WRS should work.

Startup Status

  • Wake-up of MODS1 and the LBCs is 95% complete. They could go into regular science production at any time.

  • Thanks to the photometric or nearly photometric weather, calibration and science verification of MODS1 and LBC-Blue has made excellent progress (beyond expectations). LBC-Red is a bit behind on standards.

  • Software testing and related engineering have made reasonable progress, although we have pushed off some activities into expected worse weather.

Preparations

The SX AdSec power is fixed, and the broken LBC-Red shutter coupling is fixed. Thanks to Kevin and James R respectively, and the multiple others who assisted by phone.

The plan for tonight is to observe with MODS1 on the left and LBC-red on the right. We should be able to finish the wake-up tasks for both instruments and we'll continue with photometric calibrations as clear sky and time permit.

Details

Open

01:xx Open just after sunset. Sky is mostly clear, T=12.6 degC, D=0.5 degC, wind light from SW

LBC sky flats

02:12 UT - U (LBC-B) and I (LBC-R) Skyflats

02:19 UT - LBC Red Shutter Error - stopped after 7 exposures

02:19 UT - Restarting U and I skyflat script to test for error again

02:21 UT - #7+ in Uspec flats low counts (<8k)

02:23 UT - Moving to B and I skyflats

02:27 UT - 8-10k in B skyflats, approx. 15k I skyflats

02:29 UT - Moving to V and z skyflats

02:33 UT - flat have approx. 5-10k counts

02:36 UT last flats are 023447 (Blue) and 023421 (Red)

02:39 UT - note Skyflats were not dithered - LBC was not connected to LBT

LBC wake-up (Blue and Red together)

02:41 UT - first dofpia - LBC-B and LBC-R

02:48 UT - Running Superfocus RB_rv_superfoc7.OB

02:52 UT - Running Superfocus RB_rV_superfoc7_pa180.OB

03:00 UT - Running rotatorcenter_pa0

03:02 UT - Running rotatorcenter_pa180

03:05 UT - Reconfiguring to MODS on the left - Taking LBC Bias frames at same time

MODS1 wake-up (with flat AdSec) and LBC-Red in pseudomono parallel

03:30 UT - MODS pointing check

03:46 UT - Working with MODS + LBC-R

04:09 UT - slit scan on BS9183

04:14 UT - issues with LBC-R communication - problems on power line. See Issue 5284.

04:19 UT - more issues with LBC-R - continuing with MODS - slit scan finished.

04:20 UT - doing slit scan with MODS

04:33 UT - redoing slit scan with MODS - testing tracking again

04:38 UT - Rebooting CMU (LBC)

05:17 UT - Going to dither with detxy offsets and guiding:
  • acqMODS BS9183.acq (test long slit acq and make a sum with slit that should show all stars lined up)
  • seeing is variable and not so good now... 1.2 - 1.6
   modsAlign -l mods1r.20140831.0038.fits mods1r.20140831.0039.fits
   MODS Offset Command:
   offsetxy -2.509  6.733 rel
  • confirmatory image 40 shows star a little off, but seeing was poor during acquisition.
  • second confirmatory image 41 in better seeing (now subarcsecond).
    • based on this, send offset detxy = -0.2", take 43, looks good.
    • execMODS BS9183slitscan.img does same dither as earlier +/-30, +/-15 ending with 0 again and then slit image. 2 images of 10-sec each at each position.
      • seeing about <~0.9" on guider
      • 43,44 0,0
      • 45,46 offsetxy 0 30 abs
      • 47,48 offsetxy 0 -30 abs
      • 49,50 offsetxy 0 15 abs
      • 51,52 offsetxy 0 -15 abs
      • 53,54 offsetxy 0 0 abs
    • added slit image 37 to the previous set. The +30 image appears a little to the right of the slit and the -30 one looks good.

05:47 5x Dice5 started (seeing about 1" ... going up to 0.)
    • execMODS dice10.img, execMODS -e dice10.img, ... (checkDice5 in SciPrep/Dice)
      • all these use offset type "offset abs" (i.e. radec offsets and absolute)
      • seeing is about 1" at start, creeping up gradually to 1.2 and now on the 4th out of 5 dice5's about 1.5-2", but going back down to about 1" when the megascript ended.
06:36 execMODS dice10xy.img is a single dice5 done with offset type "offsetxy rel" (note --- script did an update pointing, but the previous script did not finish at 0,0)
      • seeing about 1"
06:46 execMODS dice10rel.xy use offsetype "offset rel" (radec and relative, as for blind offsets)
      • after script ended, red fitsflush as display stopped updating at 80.
      • dice10rel

07:00 check MODS rotator centers
      • already collimated on target
      • execMODS BS9183PA0_track0 --- track preset. Star looks pretty far off center of chip. See above (updatepointing was run by the dice10xy script, and later by the dice10rel script, when star was at an offset position)
      • send an active prest to the target:
      • now track mode images at various PAs:

07:25 telescope moved to horizon so John could power cycle the LBC-Red electronics cabinet (Really only the LBC dataprobe needed to be power cycled, but it is hard to get into the cabinet on the red side, so he power cycled the cabinet again).

WRS tests with LBC-Red

The wiki log for the WRS testing is located separately on the WRS page. Instrumentation/LBCWRSSoftware

MODS1 Science and Calibration

07:31 Fe110
  • acqMODS feige110.acq
  • modsAlign -r mods1r.20140831.0097.fits
MODS Offset Command:
   offsetxy -3.728  5.266 rel
  • confirmatory image is 98, looks pretty good. Try to improve slightly by detxy offset of -0.2 dx. On 99 star appears slightly on left. 0.1 dx, 100 looks fine.
  • execMODS feige110.obs to obtain standard star spectra in all modes. airmass = 1.264. seeing 1.1-1.3"

UT 09:06 - Feige 110 - 0.6" slit dual - seeing 1" counts low on red side for standard - doing 2nd 300sec exposure

UT 09:15 - Slewing to science target FF0050 - 0.6" slit dual grating

UT 09:22 - Acquisition - seeing 0.96" Guider

UT 09:29 - Seeing 0.93" Guider - conditions are photometric

UT 10:09 - Seeing 0.84" Guider

UT 10:18 - guider image looks triangular

UT 10:38 - Seeing 0.8" Guider

UT 10:43 - 3 exposures completed, taking 1 more 1200sec exposure - Seeing 0.79" Guider

UT 10:59 - Seeing 0.89" Guider

UT 11:05 - Science object complete

UT11:10 - Fe110, now at airmass 2.020. Dual grating mode only
  • confirmatory image, 12, off to the left. x=+0.25 offset applied. 13 looks good
  • execMODS feige110dual.obs airmass 2.084, FWHM from guider 1.2 but star image stretched by atmospheric dispersion
    • red imcslock failed, but worked after a retry (I'm assuming it did not lock within the 3-min timeout window),
>> imcslock
STATUS: IMCSLOCK Blue IMCS locked-on, Red Pending BlueIMCS=LockedOn RedIMCS=Pend** ERROR: IMCSLOCK RedIMCS=Fail BlueIMCS=LockedOn
** Abort, Retry, or Ignore? > r
** Retrying command 'imcslock'
>> imcslock
STATUS: IMCSLOCK Blue IMCS locked-on, Red Pending BlueIMCS=LockedOn RedIMCS=Pending
DONE: IMCSLOCK IMCS Locked-On RedIMCS=LockedOn BlueIMCS=LockedOn
  • first one says (test, uncollimated), but not so.
  • second one - script edited & rerun, but imcslock failed again (assume timeout error). abandoning field.

UT11:34 Dither along slit with larger throw than before to confirm what looks like a slight offset between slit and detxy y-axis. Slewing to BS9107.
  • acqMODS BS9107. (lot of astigmatism to remove after moving from nearly 30 deg to 90 deg elevation - object at dec=37:20)
    >> red exptime 15
    DONE: RED EXPTIME ExpTime=15 seconds. -CLAMP EXPSTATUS=DONE
    >> red roi 3kx3k
    DONE: RED ROI ROI=(2601,5688,1,3088) EXPSTATUS=DONE
    >> slitmask in
    DONE: SLITMASK SLITMASK=5 MASKPOS=IN MASKNAME='LS5x60x0.6'
    >> red go 1
    DONE: RED IMCSLOCK Red-Channel IMCS locked-on RedIMCS=LockedOn
    >> red object BS9107 Acquisition
    DONE: RED OBJECT ImageType=OBJECT ObjectName=(BS9107 Acquisition) EXP=15.0 EXPSTATUS=INTEGRATING
    >> red exptime 15
    DONE: RED EXPTIME ExpTime=15 seconds. -CLAMP EXPSTATUS=INTEGRATING
    >> slitmask out
    DONE: SLITMASK SLITMASK=5 MASKPOS=STOW MASKNAME='LS5x60x0.6'
    >> red go 1
    ** ERROR: red GO exposure already in progress
    ** Abort, Retry, or Ignore? > i
    ** Ignoring error and continuing...
    
    Align the slit using modsAlign -l...
    
Indeed the red side was reading out when I saw this error. After it read out, I typed "i" to ignore, and the script paused but there was no red image. Had to take it manually. modsAlign -l mods1r.20140901.0015.fits mods1r.20140901.0016.fits MODS Offset Command: offsetxy -6.829 21.671 rel

Send the offset to the Telescope (Y|N)?: Y Sending offset...

  • Confirmatory image looks good (17).
  • execMODS BS9107slitscan.img will updatepointing and execute a series of field images (2x10-sec) after "offsetxy 0 dY abs" offsets.
    • about 5 minutes around transit guiding was struggling (deviations on LBTplot) and still now at 85 deg elevation.
    • images are looking elongated on-axis.

Close

12:35 closing for impending dawn. Sky is photometric.

12:45 Running biases with ROIs 2Kx2K, 3Kx3K and 1Kx1K for the dice and slitscan images taken tonight.

-- %USERSIG{JohnHill - 2014-09-01}%
Topic revision: r6 - 01 Sep 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback